Grade 12Inorganic Chemistry

How many electrons are there in one coulomb of electricity?
A
6.023×1023
B
1.64×1024
C
6.24×1018
D
6.24×1024

aman p
5 Years agoGrade 12

1 Answer

Adarsh Pal
5 Years ago

Change on 1e=1.602×1019 C

 
If total charge =1 C
 
 Total no. of e=1.602×10191=6.24×1018e
 
Hence, the correct option is (C).